# Machine Learning Guided Cooling Optimization for Data Centers

### v1

@@ Abstract
Effective data center cooling is crucial for reliable operation; however, cooling systems often exhibit inefficiencies that result in excessive energy consumption. This paper presents a three-stage, physics-guided machine learning framework for identifying and reducing cooling energy waste in high-performance computing facilities. Using one year of 10-minute resolution operational data from the Frontier exascale supercomputer, we first train a monotonicity-constrained gradient boosting surrogate that predicts facility accessory power from coolant flow rates, temperatures, and server power. The surrogate achieves a mean absolute error of 0.026 MW and predicts power usage effectiveness within $\pm$0.01 of measured values for 98.7% of test samples. In the second stage, the surrogate serves as a physics-consistent baseline to quantify excess cooling energy, revealing approximately 85 MWh of annual inefficiency concentrated in specific months, hours, and operating regimes. The third stage evaluates guardrail-constrained counterfactual adjustments to supply temperature and subloop flows, demonstrating that up to 96% of identified excess can be recovered through small, safe setpoint changes while respecting thermal limits and operational constraints. The framework yields interpretable recommendations, supports counterfactual analyses such as flow reduction during low-load periods and redistribution of thermal duty across cooling loops, and provides a practical pathway toward quantifiable reductions in accessory power. The developed framework is readily compatible with model predictive control and can be extended to other liquid-cooled data centers with different configurations and cooling requirements.
@@ I Introduction
Modern high-performance computing (HPC) systems are now among the largest single loads in data centers. A single top-tier supercomputer can draw tens of megawatts of electrical power, and almost all that power eventually leaves the facility as low-grade heat that must be removed by the cooling plant [iea2022datacenters, ebrahimi2014datacenter]. Even when an HPC facility reports an excellent Power Usage Effectiveness (PUE) of approximately 1.05, the absolute energy consumption of pumps, fans, and heat-rejection equipment remains extremely high every year [greengrid2012pue]. Small percentage improvements in cooling efficiency, therefore, translate into meaningful operating cost savings and non-trivial reductions in emissions [iea2022datacenters, gao2014machine].
@@
The Frontier supercomputer demonstrates a system that operates at high efficiency while retaining significant potential for further optimization. Frontier typically operates between 8 and 29 MW of IT (information technology) load, with an annual average of approximately 12 MW, and an observed PUE around 1.05 [frontier2022factsheet]. In other words, the facility is already running near state-of-the-art efficiency, but that still corresponds to hundreds of kilowatts of continuous accessory power for cooling and infrastructure. Our exploratory data analysis confirms that cooling overhead becomes particularly visible during low-load periods, when fixed pump and fan power is amortized over relatively little compute work.
@@
At the same time, data centers are under increasing pressure to support institutional sustainability goals and to unlock additional value from waste heat, for example, through district heating or heat pump integration [yuan2023waste, irena2020waste]. Many of the remaining opportunities are not large design changes, but micro-optimizations: slightly raising supply temperatures, slightly reducing coolant flows, or rebalancing subloops while still respecting all thermal and reliability constraints [ashrae2015thermal, doe2024bestpractices]. These are exactly the kinds of adjustments that operators are often reluctant to make manually, because the safe operating envelope is not obvious and the benefits can be smaller than day-to-day noise [gao2014machine, dechiara2020datamining]. This motivates an approach that can mine historical data for such micro-inefficiencies, propose small, safe setpoint adjustments, and quantify their expected impact before any changes are made to the live system.
@@
The study focuses on the cooling and power infrastructure serving the Frontier exascale system at Oak Ridge, which uses a warm water liquid cooling architecture with multiple secondary coolant loops that supply the compute cabinets and associated equipment [frontier2022factsheet]. In this configuration, the IT racks reject heat into three main subloops, each with its own flow and return temperature measurements, while central pumps, heat exchangers, and towers remove that heat from the facility. The electrical power drawn by this infrastructure is reported as facility accessory power and, together with compute power, underpins both the site-level PUE and our definition of cooling overhead [greengrid2012pue, frontier2022factsheet].
@@
Our analysis uses a full year of operational data from 2023, provided in the Frontier HPC & Facility Data workbook at 10-minute resolution [sun2024energy, grant2024frontier]. After cleaning and time-ordering, the dataset contains 49,869 timestamps spanning 2023-01-01 to 2023-12-31, slightly fewer than the 52,560 intervals in a complete year, reflecting scheduled downtime and short telemetry gaps. Each record includes synchronized measurements of per-loop coolant return temperatures and flow rates, overall supply temperature and flow, per-loop and total waste heat, compute power, facility accessory power, total power, and PUE, along with derived calendar fields (hour, month, weekday). Data quality of the dataset is high, with all power and flow channels are complete, and only a small fraction of overall return temperature readings are missing.
@@
Prior work on energy efficiency in HPC and data centers has followed three main tracks. The first focuses on workload and scheduling policies, aiming to reduce energy consumption while respecting performance or quality-of-service constraints. Thermal-aware and energy-aware schedulers for HPC clusters adjust job placement and dynamic voltage and frequency scaling (DVFS) settings in response to temperature and power models, often using cyber-physical formulations that couple workload, cooling, and thermal dynamics [auweter2014case, bhalachandra2015dynamic, bhalachandra2017improving]. Survey work on energy-aware scheduling emphasizes that energy is now the dominant operating cost for many HPC installations and highlights the tight coupling between temperature, cooling effort, and scheduling decisions [lazic2018datacenter]. Recent spatio-temporal algorithms explicitly model thermal diffusion in data centers, leveraging this structure to steer jobs away from hot spots and reduce cooling effort without violating thermal limits [bhalachandra2017improving].
@@
The second major line of work focuses on cooling and plant aspects, employing advanced control to reduce the power of chiller, pump, and fan systems. Model-predictive control (MPC) has been successfully applied to data center chiller plants, using physics-based models and forecasts of load and weather to optimize water temperatures, flow rates, and equipment staging under operational constraints [wang2022physics]. More recently, deep reinforcement learning (DRL) has been used to learn cooling policies directly from data, including algorithms that co-optimize IT and cooling systems [moriyama2018reinforcement], transform cooling optimization into a DRL control problem [wang2025tropical], jointly optimize job scheduling and cooling [meta2024simulator], and operate free-cooled data centers in tropical climates [beloglazov2011energy]. These approaches demonstrate substantial energy savings, but they typically rely on complex black-box policies and are often validated on air-cooled facilities rather than warm water liquid-cooled supercomputers.
@@
A third, emerging strand combines safety-aware learning with physics-guidance, recognizing that purely black-box controllers are difficult to deploy in critical infrastructure. Recent work on safe RL for building and plant control introduces explicit constraints, action-space pruning, and physics-inspired monotonicity checks to keep learned policies within safe operating envelopes [wang2022physics, wang2025tropical]. However, most of these studies either operate at the building-HVAC level or treat the data center as a single thermal zone; they rarely expose the fine-grained liquid-cooling loop behavior or accessory-power breakdown needed to reason about micro-optimizations such as small supply-temperature shifts or loop-specific flow trims. To our knowledge, the publicly released Frontier HPC & facility dataset [sun2024energy, grant2024frontier] has not yet been studied based on physics-guided, guardrail-constrained micro-tuning of liquid-cooling operation at 10-minute resolution.
@@ II Methodology
In the third stage, the surrogate from Step 1 is used as a counterfactual oracle. At each 10-minute interval, we estimate the accessory-power reduction that would have been achievable under a minimally perturbed cooling action, subject to the same safety constraints. To keep the policy interpretable and operationally realistic, we restrict attention to two familiar control knobs: the coolant supply temperature $T_{\text{sup}}$, and the per-loop coolant flows $Q_{1},Q_{2},Q_{3}$.
@@ III-A Surrogate performance
The physics-guided LightGBM surrogate reproduces Frontier’s accessory power $P_{\mathrm{acc}}$ with errors on the order of only a few tens of kilowatts. On a random 80/20 split of the 2023 dataset, the test-set MAE is $0.0259~\mathrm{MW}$ and the RMSE is $0.0387~\mathrm{MW}$, with $R^{2}=0.79$ and $\mathrm{SMAPE}\approx 4.2\%$. The weighted absolute percentage error (WAPE) is $4.3\%$, and the log-scaled error (RMSLE) is $0.0233$. Training errors are slightly lower (MAE $0.0193~\mathrm{MW}$, RMSE $0.0293~\mathrm{MW}$, $R^{2}=0.88$), indicating a modest but expected generalization gap without evidence of severe overfitting. When mapped into PUE, the test-set error remains very small ($\mathrm{PUE\_MAE}=0.00225$, $\mathrm{PUE\_RMSE}=0.00406$), with $98.7\%$ of test points within $\pm 0.01$ of the measured PUE (Table I).
[…]
TABLE I: Performance of the physics-guided LightGBM surrogate for accessory power $P_{\mathrm{acc}}$ on a random 80/20 train-test split.
@@ III-B Excess characterization
Using the Step 1 surrogate as a physics-consistent baseline, we quantify excess accessory power at 10-minute resolution over the 2023 Frontier dataset and then aggregate it to energy and cost. Over the full year, the implied excess cooling energy is $85.2~\mathrm{MWh}$, corresponding to an energy cost of approximately \$5,100 at \$60/MWh and a small additional demand charge, for a total annual excess cost of about \$5,100 (Table II). Although this represents only a fraction of the overall cooling energy, it is a non-negligible resource that can be targeted by micro-optimizations.
TABLE II: Annual excess cooling energy and cost relative to the surrogate baseline.
@@ IV Conclusion
Step 3 then uses the same surrogate and feature space as a counterfactual oracle, exploring small, guardrail-constrained adjustments to supply temperature and subloop flows at each 10-minute interval. Under pure physics guardrails, the counterfactual policy can recover roughly $127~\mathrm{MWh}$ of cooling energy; when capped by the Step 2 excess on a per-interval basis, this shrinks to $\sim 82~\mathrm{MWh}$, or about $96\%$ of the identified excess. After applying conservative reviewer filters (in-distribution checks, materiality thresholds, and simple hysteresis), the resulting sparse policy touches only a few percent of time steps yet still delivers about $13~\mathrm{MWh}$ per year of credible savings through micro-adjustments at the tens-of-kilowatt level. Together, these results suggest that even in a facility with already excellent PUE, there is a measurable band of recoverable cooling energy that can be accessed via safe, interpretable tuning of existing setpoints rather than major hardware changes. Together, these results demonstrate that safe, interpretable micro-adjustments can recover up to 96% of identified cooling inefficiencies, corresponding to approximately 82 MWh annually under the Step 2 excess baseline, with a more conservative estimate of 13–15 MWh under strict reviewer filters.
The present study is subject to several limitations that motivate future work. First, all results are based on offline counterfactual evaluation of a single year of historical data from one installation; the surrogate and policy are not yet validated under prospectively collected telemetry or in closed-loop operation, and the conclusions may not transfer directly to other liquid-cooled sites with different hardware, controls, or climates. Second, the action space is deliberately narrow (small $\Delta T_{\mathrm{sup}}$ and modest flow trims), and the guardrails are conservative, so the reported savings should be viewed as lower bounds relative to what more aggressive but still safe policies might achieve. Third, the framework currently treats IT workload as an exogenous input and does not co-optimize scheduling and cooling, nor does it model interactions with upstream plant equipment (e.g., chillers, towers, or heat-recovery systems).

### v2

# Machine Learning Guided Cooling System Optimization for Data Center

@@ Abstract
Effective data center cooling is crucial for reliable operation; however, cooling systems often exhibit inefficiencies that result in excessive energy consumption. This paper presents a three-stage, physics-guided machine learning framework for identifying and reducing cooling energy waste in high-performance computing facilities. Using one year of 10-minute resolution operational data from the Frontier exascale supercomputer, we first train a monotonicity-constrained gradient boosting surrogate that predicts facility accessory power from coolant flow rates, temperatures, and server power. The surrogate achieves a mean absolute error of 0.026 MW and predicts power usage effectiveness within $\pm$0.01 of measured values for 98.7% of test samples. In the second stage, the surrogate serves as a physics-consistent baseline to quantify excess cooling energy, revealing approximately 85 MWh of annual inefficiency concentrated in specific months, hours, and operating regimes. The third stage evaluates guardrail-constrained counterfactual adjustments to supply temperature and subloop flows, demonstrating that up to 96% of identified excess can be recovered through small, safe setpoint changes while respecting thermal limits and operational constraints. The framework yields interpretable recommendations, supports counterfactual analyses such as flow reduction during low-load periods and redistribution of thermal duty across cooling loops, and provides a practical pathway toward quantifiable reductions in accessory power. The developed framework is readily compatible with model predictive control and provides a template that, with site-specific recalibration, could be adapted to other liquid-cooled data centers with different configurations and cooling requirements.
@@ I Introduction
Modern high-performance computing (HPC) systems are now among the largest single loads in data centers. A single top-tier supercomputer can draw tens of megawatts of electrical power, and almost all that power eventually leaves the facility as low-grade heat that must be removed by the cooling plant [10, 7]. Even when an HPC facility reports an excellent Power Usage Effectiveness (PUE) of approximately 1.05, the absolute energy consumption of pumps, fans, and heat-rejection equipment remains extremely high every year [17]. Small percentage improvements in cooling efficiency, therefore, translate into meaningful operating cost savings and non-trivial reductions in emissions [10, 8].
@@
The Frontier supercomputer demonstrates a system that operates at high efficiency while retaining significant potential for further optimization. Frontier typically operates between 8 and 29 MW of IT (information technology) load, with an annual average of approximately 12 MW, and an observed PUE around 1.05 [15]. In other words, the facility is already running near state-of-the-art efficiency, but that still corresponds to hundreds of kilowatts of continuous accessory power for cooling and infrastructure. Our exploratory data analysis confirms that cooling overhead becomes particularly visible during low-load periods, when fixed pump and fan power is amortized over relatively little compute work.
@@
At the same time, data centers are under increasing pressure to support institutional sustainability goals and to unlock additional value from waste heat, for example, through district heating or heat pump integration [21, 11]. Many of the remaining opportunities are not large design changes, but micro-optimizations: slightly raising supply temperatures, slightly reducing coolant flows, or rebalancing subloops while still respecting all thermal and reliability constraints [1, 18]. These are exactly the kinds of adjustments that operators are often reluctant to make manually, because the safe operating envelope is not obvious and the benefits can be smaller than day-to-day noise [8, 6]. This motivates an approach that can mine historical data for such micro-inefficiencies, propose small, safe setpoint adjustments, and quantify their expected impact before any changes are made to the live system.
@@
The study focuses on the cooling and power infrastructure serving the Frontier exascale system at Oak Ridge, which uses a warm water liquid cooling architecture with multiple secondary coolant loops that supply the compute cabinets and associated equipment [15]. In this configuration, the IT racks reject heat into three main subloops, each with its own flow and return temperature measurements, while central pumps, heat exchangers, and towers remove that heat from the facility. The electrical power drawn by this infrastructure is reported as facility accessory power and, together with compute power, underpins both the site-level PUE and our definition of cooling overhead [17, 15].
@@
Our analysis uses a full year of operational data from 2023, provided in the Frontier HPC & Facility Data workbook at 10-minute resolution [16, 9]. After cleaning and time-ordering, the dataset contains 49,869 timestamps spanning 2023-01-01 to 2023-12-31, slightly fewer than the 52,560 intervals in a complete year, reflecting scheduled downtime and short telemetry gaps. Each record includes synchronized measurements of per-loop coolant return temperatures and flow rates, overall supply temperature and flow, per-loop and total waste heat, compute power, facility accessory power, total power, and PUE, along with derived calendar fields (hour, month, weekday). Data quality of the dataset is high, with all power and flow channels are complete, and only a small fraction of overall return temperature readings are missing. Table I summarizes the key variables and their characteristics in the Frontier dataset.
TABLE I: Overview of the Frontier HPC & Facility Dataset.
Variable Description Typical Range
$P_{\text{IT}}$ IT/compute power 8–29 MW
$P_{\text{acc}}$ Facility accessory power 0.5–1.1 MW
$T_{\text{sup}}$ Coolant supply temperature 18–25C
$T_{r,i}$ Subloop return temperatures 25–40C
$Q_{i}$ Subloop coolant flow rates Variable
$Q_{\text{heat}}$ Total waste heat 5–25 MW
PUE Power usage effectiveness 1.03–1.10
Resolution: 10 minutes; Period: 2023-01-01 to 2023-12-31
Total records: 49,869 (after cleaning)
@@
Prior work on energy efficiency in HPC and data centers has followed three main tracks. The first focuses on workload and scheduling policies, aiming to reduce energy consumption while respecting performance or quality-of-service constraints. Thermal-aware and energy-aware schedulers for HPC clusters adjust job placement and dynamic voltage and frequency scaling (DVFS) settings in response to temperature and power models, often using cyber-physical formulations that couple workload, cooling, and thermal dynamics [2, 4, 5]. Survey work on energy-aware scheduling emphasizes that energy is now the dominant operating cost for many HPC installations and highlights the tight coupling between temperature, cooling effort, and scheduling decisions [12]. Recent spatio-temporal scheduling algorithms, which explicitly model thermal diffusion across the physical layout of data centers, leverage this structure to steer jobs away from hot spots and reduce cooling effort without violating thermal limits [5].
@@
The second major line of work focuses on cooling and plant aspects, employing advanced control to reduce the power of chiller, pump, and fan systems. Model-predictive control (MPC) has been successfully applied to data center chiller plants, using physics-based models and forecasts of load and weather to optimize water temperatures, flow rates, and equipment staging under operational constraints [19]. More recently, deep reinforcement learning (DRL) has been used to learn cooling policies directly from data, including algorithms that co-optimize IT and cooling systems [14], transform cooling optimization into a DRL control problem [20], jointly optimize job scheduling and cooling [13], and operate free-cooled data centers in tropical climates [3]. These approaches demonstrate substantial energy savings, but they typically rely on complex black-box policies and are often validated on air-cooled facilities rather than warm water liquid-cooled supercomputers.
@@
A third, emerging strand combines safety-aware learning with physics-guidance, recognizing that purely black-box controllers are difficult to deploy in critical infrastructure. Recent work on safe RL for building and plant control introduces explicit constraints, action-space pruning, and physics-inspired monotonicity checks to keep learned policies within safe operating envelopes [19, 20]. Monotonicity constraints are particularly valuable because they encode domain knowledge that certain relationships should only increase or decrease: for example, higher coolant flow or greater heat load should never reduce cooling power, and enforcing this prevents the model from learning spurious correlations that could lead to unsafe recommendations. However, most of these studies either operate at the building-HVAC level or treat the data center as a single thermal zone; they rarely expose the fine-grained liquid-cooling loop behavior or accessory-power breakdown needed to reason about micro-optimizations such as small supply-temperature shifts or loop-specific flow trims. To our knowledge, the publicly released Frontier HPC & facility dataset [16, 9] has not yet been studied based on physics-guided, guardrail-constrained micro-tuning of liquid-cooling operation at 10-minute resolution.
@@ II Methodology
We selected Light Gradient Boosting Machine (LightGBM) as the surrogate model for several reasons. First, gradient boosting methods handle mixed feature types (continuous, categorical, temporal) naturally without extensive preprocessing. Second, LightGBM natively supports monotonicity constraints on selected features, which is essential for encoding physics-based priors. Third, compared to deep learning approaches, LightGBM provides faster training, lower computational overhead, and more interpretable feature importance rankings, all of which are valuable for operational deployment in critical infrastructure. Finally, tree-based ensembles have demonstrated strong performance on tabular regression tasks with moderate dataset sizes, making them well-suited to this 10-minute resolution telemetry problem.
@@
In the third stage, the surrogate from Step 1 is used as a counterfactual oracle. The core idea is to ask: “What if we had operated with slightly different setpoints at each historical timestamp?” By replaying the year with small, hypothetical adjustments to supply temperature and flow rates, we can estimate the energy that could have been saved without actually modifying the live system. The guardrails ensure that these hypothetical adjustments remain physically plausible (e.g., heat must still be removed) and operationally safe (e.g., temperatures and flows stay within proven bounds). At each 10-minute interval, we estimat the accessory-power reduction that would have been achievable under a minimally perturbed cooling action, subject to the same safety constraints. To keep the policy interpretable and operationally realistic, we restrict attention to two familiar control knobs: the coolant supply temperature $T_{\text{sup}}$, and the per-loop coolant flows $Q_{1},Q_{2},Q_{3}$.
@@ III-A Surrogate performance
The physics-guided LightGBM surrogate reproduces Frontier’s accessory power $P_{\mathrm{acc}}$ with errors on the order of only a few tens of kilowatts. On a random 80/20 split of the 2023 dataset, the test-set MAE is $0.0259~\mathrm{MW}$ and the RMSE is $0.0387~\mathrm{MW}$, with $R^{2}=0.79$ and $\mathrm{SMAPE}\approx 4.2\%$. The weighted absolute percentage error (WAPE) is $4.3\%$, and the log-scaled error (RMSLE) is $0.0233$. Training errors are slightly lower (MAE $0.0193~\mathrm{MW}$, RMSE $0.0293~\mathrm{MW}$, $R^{2}=0.88$), indicating a modest but expected generalization gap without evidence of severe overfitting. When mapped into PUE, the test-set error remains very small ($\mathrm{PUE\_MAE}=0.00225$, $\mathrm{PUE\_RMSE}=0.00406$), with $98.7\%$ of test points within $\pm 0.01$ of the measured PUE (Table II).
TABLE II: Performance of the physics-guided LightGBM surrogate for accessory power $P_{\mathrm{acc}}$ on a random 80/20 train-test split.
@@ III-B Excess characterization
[…]
Using the Step 1 surrogate as a physics-consistent baseline, we quantify excess accessory power at 10-minute resolution over the 2023 Frontier dataset and then aggregate it to energy and cost. Over the full year, the implied excess cooling energy is $85.2~\mathrm{MWh}$, corresponding to an energy cost of approximately \$5,100 at \$60/MWh and a small additional demand charge, for a total annual excess cost of about \$5,100 (Table III). Although this represents only a fraction of the overall cooling energy, it is a non-negligible resource that can be targeted by micro-optimizations.
TABLE III: Annual excess cooling energy and cost relative to the surrogate baseline.
@@ III-C Counterfactual savings and safety
To assess the sensitivity of these estimates to surrogate uncertainty, we note that the Step 1 test MAE of 0.026 MW corresponds to an annual energy uncertainty of approximately $\pm 13$ MWh if the error were systematic across all intervals. However, because residuals are approximately symmetric and centered near zero (mean $\approx 0.0007$ MW), random errors largely cancel over aggregation. The capped savings of 82.1 MWh exceed this uncertainty band by a factor of six, providing confidence that the reported savings are not artifacts of model error. For the more conservative reviewer-pass estimate of 13.4 MWh, the margin is smaller, and these savings should be interpreted as order-of-magnitude guidance rather than precise predictions.
@@ IV Conclusion
Step 3 then uses the same surrogate and feature space as a counterfactual oracle, exploring small, guardrail-constrained adjustments to supply temperature and subloop flows at each 10-minute interval. Under pure physics guardrails, the counterfactual policy can recover roughly $127~\mathrm{MWh}$ of cooling energy; when capped by the Step 2 excess on a per-interval basis, this shrinks to $\sim 82~\mathrm{MWh}$, or about $96\%$ of the identified excess. After applying conservative reviewer filters (in-distribution checks, materiality thresholds, and simple hysteresis), the resulting sparse policy touches only a few percent of time steps yet still delivers about $13~\mathrm{MWh}$ per year of credible savings through micro-adjustments at the tens-of-kilowatt level. Together, these results demonstrate that safe, interpretable micro-adjustments can recover up to 96% of identified cooling inefficiencies, corresponding to approximately 82 MWh annually under the Step 2 excess baseline, with a more conservative estimate of 13–15 MWh under strict reviewer filters. These findings suggest that even in a facility with already excellent PUE, there remains a measurable band of recoverable cooling energy accessible through physics-guided setpoint tuning.
The present study is subject to several limitations that motivate future work. First, all results are based on offline counterfactual evaluation of a single year of historical data from one installation; the surrogate and policy are not yet validated under prospectively collected telemetry or in closed-loop operation, and the conclusions would require site-specific recalibration before being applied to other liquid-cooled sites with different hardware, controls, or climates. Second, the action space is deliberately narrow (small $\Delta T_{\mathrm{sup}}$ and modest flow trims), and the guardrails are conservative, so the reported savings should be viewed as lower bounds relative to what more aggressive but still safe policies might achieve. Third, the framework currently treats IT workload as an exogenous input and does not co-optimize scheduling and cooling, nor does it model interactions with upstream plant equipment (e.g., chillers, towers, or heat-recovery systems).
@@
V Code Availability
The code used in this study can be accessed at https://github.com/m-iml/ML-Optimization-Data-Centers.